OpenStudy (anonymous):

Explain the identity property and how it is used. What are the identity elements for addition and multiplication?

OpenStudy (anonymous):

M*a=a*M = M , where a= multiplication identity element

OpenStudy (anonymous):

5*1 = 1*5 =5

OpenStudy (anonymous):

also for the addition, we consider zero as the identity addition element ==> M+0 = 0+M = M

OpenStudy (sasogeek):

The identity property is that which returns the value of an operation as exactly one of the objects involved.... the id element for addition in 0 and that of multiplication is 1 eg. if i have 1*5, my result is 5. 5 was an object in the operation but was also the value returned. 5+0=5..... i hope u can get understand my point

OpenStudy (sasogeek):

mmm i meant to say i hope you get my point lol :P